OpenStudy (anonymous):

how do you solve tan^-1 (x-(1/x) / 2)

OpenStudy (helder_edwin):

\[ \large \tan^{-1}(x-\frac{\frac{1}{x}}{2})=? \]

OpenStudy (anonymous):

nono ill draw it right now on a tablet so you could see much more clear!

OpenStudy (helder_edwin):

ok

OpenStudy (anonymous):

@helder_edwin

jimthompson5910 (jim_thompson5910):

tan(angle) = opposite/adjacent

jimthompson5910 (jim_thompson5910):

tan(theta) = opp/adj tan(theta) = (x-1/x)/2 theta = tan^(-1) ( (x-1/x)/2 )

OpenStudy (helder_edwin):

ok. but is it an equation?

jimthompson5910 (jim_thompson5910):

so we basically have something like this 

jimthompson5910 (jim_thompson5910):

hmm seems like something is missing

OpenStudy (anonymous):

okay i will give you the whole question!

OpenStudy (anonymous):

the question is to verify the trig identity of this! @jim_thompson5910

OpenStudy (anonymous):